Consider the Mungo River where it is 200

meters wide and has a uniform flow. A small

boat sailing at full speed takes 5 minutes to

cover 100 meters when sailing upstream, but

only 2 minutes when sailing downstream. What

is the shortest time in which the boat can cross

the Mungo to the nearest point on the opposite

bank?


1
Expert's answer
2021-02-17T11:01:06-0500

"\begin{cases}
 5=\frac{100}{v_b-v_r},\\
 2=\frac{100}{v_b+v_r};
\end{cases}"

"\begin{cases}
 {v_b-v_r}=20,\\
 {v_b+v_r}=50;
\end{cases}"

"\begin{cases}
 v_b=35~\frac{m}{min},\\
 {v_r}=15~\frac{m}{min};
\end{cases}"

"t=\frac{200}{\sqrt{v_b^2+v_r^2}}"

"t\approx 5.25 ~min."